# CounterQuill: Investigating the Potential of Human-AI Collaboration in Online Counterspeech Writing

### v1

@@ 8. CONCLUSION
    Have you used Large Language Model (LLM) tools or an AI chatbot, such as ChatGPT, Google Gemini or Claude?
    1. (a)
      Yes
    2. (b)
      No
@@
    Have you ever countered hate content on online social media?
    1. (a)
      Yes
    2. (b)
      No
@@
    1. (a)
      Woman
    2. (b)
      Man
    3. (c)
      Trans man
    4. (d)
      Trans woman
    5. (e)
      Genderqueer
    6. (f)
      Agender
    7. (g)
      Autigender
    8. (h)
      Prefer not to answer
    9. (i)
      Other
@@
    1. (a)
      Asian/Asian American
    2. (b)
      Black/African American
    3. (c)
      Hispanic/Latina/o/x
    4. (d)
      Middle Eastern/North African
    5. (e)
      Native American/American Indian
    6. (f)
      Native Hawaiian/Pacific Islander
    7. (g)
      White/European American
    8. (h)
      Other
@@
    1. (a)
      Less than high school
    2. (b)
      High school graduate
    3. (c)
      Some college
    4. (d)
      Associate degree
    5. (e)
      Bachelor’s degree
    6. (f)
      Some postgraduate work
    7. (g)
      Master’s degree
    8. (h)
      Doctorate degree
@@
    1. (a)
      Very conservative
    2. (b)
      Conservative
    3. (c)
      Moderate
    4. (d)
      Liberal
    5. (e)
      Very Liberal
  11. (11)
    We would like to invite you to a follow-up interview study based on your responses. Are you interested in participating?
    1. (a)
      Yes
    2. (b)
      No
  12. (12)
    Are you comfortable speaking and participating in this study in English?
    1. (a)
      Yes
    2. (b)
      No
  13. (13)
    Do you have a laptop/PC?
    1. (a)
      Yes
    2. (b)
      No
@@ Appendix B Guiding interview questions
Appendix C post-study questionnaire
@@
    Very Low
     1  2  3  4  5  6  7  Very High
@@
    Strongly Disagree
     1  2  3  4  5  6  7  Strongly Agree
  3. (3)
    How confident do you feel about the quality of the counter-speech you wrote with the help of ChatGPT (condition A) / CounterQuill (condition B)?
    Strongly Disagree
     1  2  3  4  5  6  7  Strongly Agree
  4. (4)
    To what extent did ChatGPT (condition A) / CounterQuill (condition B) provide customized assistance that was helpful for your specific counter-speech writing needs?
    Not at all
     1  2  3  4  5  6  7  To a great extent
@@
    Not at all likely
     1  2  3  4  5  6  7  Extremely likely
Appendix D quiz questions
@@
Appendix E CounterQuill prompt
Because CounterQuill prompts are long, we created a GitHub repository to save our prompts. See https://github.com/AnonymousUseOnly/Group25 or Supplementary Materials for details.
Appendix F Examples of Hate Speech
Warning: Table A1 lists examples of hate speech used by user study in order of participants. The table contains speech examples that may be offensive or upsetting. These examples are presented solely for academic and research purposes and do not reflect the views of the author.
@@ Appendix F Examples of Hate Speech
#sexual_orient Lesbians are often stereotyped as ‘tomboys,’ rejecting traditional female roles and appearances, and when I see them, I get uncomfortable because they’re not feminine at all.
@@
Appendix G Results of Welch’s Two-Sample t-tests
We conducted Welch’s Two-Sample t-tests that accounts for unequal variances between groups (Sakai, 2016), as detailed in Tables A2 and A3. The results confirmed that the sequence—whether participants began with ChatGPT or CounterQuill—did not significantly influence their performance or feedback on the system.
Table A2. Results of Welch’s Two-Sample t-tests assessing the effect of counterbalancing the order of system usage (ChatGPT First vs. CounterQuill First) on NASA-TLX scores. The analysis was conducted to determine if the sequence in which participants used ChatGPT and CounterQuill influenced their perceived workload across various dimensions (Mental Demand, Physical Demand, Temporal Demand, Performance, Effort, and Frustration). No significant differences were found (all p ¿ 0.05), indicating that the order of system usage did not affect participants’ performance or feedback.
Platform NASA-TLX Items $Mean\pm SD$ t-value p-value 95% CI
ChatGPT First CounterQuill First
Performance 4.00$\pm$1.05 3.80$\pm$1.14 0.408 0.688 [-0.830, 1.230]
Effort 2.60$\pm$1.17 2.20$\pm$0.92 0.849 0.408 [-0.594, 1.394]
Frustration 3.40$\pm$1.90 2.60$\pm$2.17 0.878 0.392 [-1.118, 2.718]
Performance 6.20$\pm$0.79 6.50$\pm$0.71 -0.896 0.382 [-1.004, 0.404]
Effort 4.20$\pm$1.55 3.90$\pm$1.79 0.400 0.694 [-1.276, 1.876]
Frustration 1.70$\pm$0.82 1.30$\pm$0.67 1.188 0.251 [-0.309, 1.109]
Table A3. Results of Welch’s Two-Sample t-tests assessing the impact of counterbalancing the order of system usage (ChatGPT First vs. CounterQuill First) on responses to customized questions. The questions evaluated various aspects including Hate Speech (HS) Identification Confidence, Brainstorming Effectiveness, Self-Efficacy in Counterspeech (CS) Writing, Engagement with AI, Satisfaction with Counterspeech, and Willingness to Post Online. The analysis showed no significant differences in responses based on the order of system usage (all p ¿ 0.05), suggesting that the sequence in which participants used ChatGPT and CounterQuill did not influence their responses to these customized questions.
Platform Category of Question $Mean\pm SD$ $t$-value $p$-value 95% CI
ChatGPT First CounterQuill First

### v2

@@ Appendix A Preliminary survey questions
    Have you used Large Language Model (LLM) tools or an AI chatbot, such as ChatGPT, Google Gemini or Claude? (Yes, No)
@@
    Have you ever countered hate content on online social media? (Yes, No)
@@
  5. (5)
    If you were writing a counterspeech on social media, would you use AI technology like ChatGPT to help you? (1 = Definitely not, 5 = Definitely yes)
  6. (6)
    Why would or wouldn’t you use AI technology to help you write a counterspeech?
@@ Appendix B Guiding interview questions
Appendix C Post-Study NASA Question
@@
Appendix D Post-Study Custom Question
@@
Appendix E Quiz questions
@@
Appendix F CounterQuill prompt
Because CounterQuill prompts are long, we created a GitHub repository to save our prompts. See https://github.com/AnonymousUseOnly/CounterQuill_Prompt for details.
Appendix G Results of Welch’s Two-Sample t-tests
We conducted Welch’s Two-Sample t-tests that accounts for unequal variances between groups (Sakai, 2016), as detailed in Tables A1 and A2. The results confirmed that the sequence—whether participants began with ChatGPT or CounterQuill—did not significantly influence their performance or feedback on the system.
Table A1. Welch’s Two-Sample t-tests comparing the effect of system usage order (ChatGPT First vs. CounterQuill First) on NASA-TLX workload dimensions (Mental, Physical, Temporal Demands, Performance, Effort, and Frustration). No significant differences were found (all p ¿ 0.05), indicating order did not affect participants’ performance or feedback.
Platform NASA-TLX Items $Mean\pm SD$ t-value p-value 95% CI
ChatGPT First CounterQuill First
Performance 4.00$\pm$1.05 3.80$\pm$1.14 0.408 0.688 [-0.830, 1.230]
Effort 2.60$\pm$1.17 2.20$\pm$0.92 0.849 0.408 [-0.594, 1.394]
Frustration 3.40$\pm$1.90 2.60$\pm$2.17 0.878 0.392 [-1.118, 2.718]
Performance 6.20$\pm$0.79 6.50$\pm$0.71 -0.896 0.382 [-1.004, 0.404]
Effort 4.20$\pm$1.55 3.90$\pm$1.79 0.400 0.694 [-1.276, 1.876]
Frustration 1.70$\pm$0.82 1.30$\pm$0.67 1.188 0.251 [-0.309, 1.109]
Table A2. Welch’s Two-Sample t-tests comparing the effect of system usage order (ChatGPT First vs. CounterQuill First) on responses to customized questions measuring HS Identification Confidence, Brainstorming Effectiveness, Self-Efficacy in CS Writing, Engagement with AI, Satisfaction with Counterspeech, and Willingness to Post Online. No significant differences were found (all p ¿ 0.05), indicating order did not influence participants’ responses.
Platform Category of Question $Mean\pm SD$ $t$-value $p$-value 95% CI
ChatGPT First CounterQuill First
Appendix H Examples of hate speech
Warning: Table LABEL:speech_examples_table contains potentially offensive or upsetting hate speech examples used in the user study. These are presented for academic and research purposes only and do not reflect the author’s views.
@@ Appendix H Examples of hate speech
#sexual_orient Lesbians are often stereotyped as ’tomboys,’ rejecting traditional female roles and appearances, and when I see them, I get uncomfortable because they’re not feminine at all.